2x + y = 1
4x + 2y = -1

Answer by swincher4391(1107)

2x + y = 1
4x + 2y = -1
use elimination.
multiply the top equation by 2 then subtract the resulting equations

4x + 2y = 2
4x + 2y = -1
--------------
0 + 0 = 3
0=3 is a false statement so the answer is NO SOLUTION